OpenStudy (anonymous):

what is the solution to sinx = sin3x

OpenStudy (dumbcow):

same solution as x = 3x x = 0 Note for any multiple of pi, the sin is also 0 solution is 0, pi, 2pi ... npi

OpenStudy (dumbcow):

wait that is only half the solution, sorry to get full solution you need expand sin(3x) http://en.wikipedia.org/wiki/Trig_identity#Double-angle.2C_triple-angle.2C_and_half-angle_formulae now you have \[\sin x = -4\sin^3 x +3\sin x\] which factors to \[2 \sin x(2 \sin^2 x -1) = 0\] which gives solutions \[x = n \pi , \frac{\pi}{4} + n \frac{\pi}{2}\]

OpenStudy (anonymous):

Alternatively use \[\sin A -\sin B =2\cos (\frac{ A+B }{2 })\sin (\frac{ A-B }{2 })\]

OpenStudy (amoodarya):

\[\sin3x=sinx \\3x=x+2k \pi\\3x=\pi-x+2k \pi\\3x=x+2k \pi \rightarrow 2x=2k \pi \rightarrow x=k \pi\\3x=\pi-x+2k \pi \rightarrow 4x=\pi+2k \pi \rightarrow x=\frac{ \pi }{ 4}+\frac{ k \pi }{ 2 }\]

OpenStudy (amoodarya):

\[k \in Z\]

OpenStudy (amoodarya):